# Effects of Near-Field Hydrodynamic Interactions on Bacterial Dynamics Near a Solid Surface

### v1

@@ III.1 Simulation Without Brownian Motion
Figure 3: The normalized velocities of a bacterium in the approach stage. (a) The vertical translational velocity $U_{\perp}$ as a function of reduced height $h/R_{b}$ normalized by the bulk velocity $U_{\infty}$. (b) The vertical rotational velocity $W_{\perp}$ as a function of reduced height $h/R_{b}$ normalized by the bulk velocity $W_{\infty}$.
@@
We first examine the effect of near-field HIs on the bacterial mobility near a no-slip solid surface, excluding Brownian motion. During this process, only hydrodynamic and DLVO interactions are considered. Surface entrapment is typically divided into three sequential stages: approach, reorientation, and surface swimming Bianchi, Saglimbeni, and Di Leonardo (2017); Bianchi et al. (2019); Liu, Jin, and An (2025). The translational velocity of bacteria in the bulk fluid is about $U_{\infty}=14.5$ $\mu\mathrm{m}\mathrm{/}\mathrm{s}$ for the parameters provided in Table 1, consistent with the results of the experimental measurement Molaei et al. (2014). In Fig. 3, we plot the normalized vertical translational $U_{\perp}/U_{\infty}$ and rotational velocities $W_{\perp}/W_{\infty}$ as a function of the reduced height $h/R_{b}$. It is evident that the resistance function $Y_{\perp}^{A}$ increases as the separation distance $h$ decreases and the vertical velocity in the approach stage decreases accordingly. The function $Y_{\perp}^{A}$ in Eq. 23 diverges to infinity as the height $h$ approaches zero, resulting in the translational velocity of the bacteria approaching zero. Consequently, the bacterial vertical motion is significantly restricted, and they are effectively confined to the surface. However, the DLVO force prevents the bacteria from approaching the surface infinitely. Instead, the separation distance between the cell body and the surface is on the order of nanometers Vigeant and Ford (1997); Liu, Jin, and An (2025), which causes bacteria to move in circular trajectories on the surface.
Similarly, resistance functions $Y_{\parallel}^{A}$, $Y_{\parallel}^{C}$, and $Y_{\perp}^{C}$ increase as the reduced height $h/R_{b}$ decreases (Figs. 2(a) and (c)), reducing the rotational velocity of the bacteria. In the absence of a surface, the bacterial rotational velocity is about $W_{\infty}=18.3$ $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$, based on the parameters in Table 1. As bacteria approach the surface, the rotation of the cell body is suppressed, while flagellar rotation is accelerated. The vertical rotational velocity experiences a slight increase before decreasing, as depicted in Fig. 3(b). However, the closer the bacteria are to the surface, the greater the suppression of their translational velocity. The net bacterial velocity is the result of the competition between translational and rotational velocities. Due to the more significant suppression of translational velocity relative to rotational velocity, vertical translational velocity monotonically decreases with decreasing height $h$, as shown in Fig. 3(a).
@@
Figure 4: Bacterial velocities and radius of curvature of their circular motion in the surface swimming stage. (a) Horizontal translational velocity $U_{\parallel}$ as a function of the stable height $h^{*}$. (b) Vertical rotational velocity $W_{\perp}$ as a function of the stable height $h^{*}$. (c) Radius of curvature, $r$, as a function of the stable height $h^{*}$. (d) Radius of curvature, $r$, as a function of the horizontal translational velocity $U_{\parallel}$.
@@
On a no-slip solid surface, there exists a stable fixed point $\{h^{*},\Psi^{*}\}$ in the phase plane $\{h,\Psi\}$ Shum, Gaffney, and Smith (2010); Shum and Gaffney (2015), where the bacterium performs stable circular motions on the surface. Experimental studies have revealed that the radius of curvature is dependent on the bacterial swimming speed Liu et al. (2024). The horizontal translational and vertical rotational velocities of bacteria are solely determined by this stable fixed point $\{h^{*},\Psi^{*}\}$, with fixed bacterial morphology and motor rotation rate. Bacteria swim in circular motion near a solid surface during the surface swimming stage Lauga et al. (2006). The radius of curvature, $r$, of the circular motion is given by Liu, Jin, and An (2025)
@@
where $\Psi^{*}$ is the stable inclination angle, $\phi$ is the azimuth angle of the rotational velocity in spherical coordinates, $U_{\parallel}=\sqrt{U_{x}^{2}+U_{y}^{2}}$ is the translational velocity parallel to the surface and $W_{\perp}=|W_{z}-\left(W_{x}\cos\phi+W_{y}\sin\phi\right)\cot\Psi^{*}|$ is the rotational velocity perpendicular to the surface (along the $z$-axis).
@@
Observations from Eqs. 22-26 reveal that individual resistance functions of the cell body exhibit varying dependencies on the separation distance $h$. Figs. 4(a)-(b) show that the horizontal translational velocity $U_{\parallel}$ increases with stable height $h^{*}$, while the vertical rotational velocity $W_{\perp}$ decreases as $h^{*}$ increases. As shown in Fig. 4(c), the radius of curvature of the bacterial circular motion increases monotonically with stable height. For stable heights that exceed $40$ $\mathrm{n}\mathrm{m}$, the radius of curvature exhibits a linear relationship with $h^{*}$ Liu et al. (2024). As shown in Fig. 4(d), the radius of curvature also increases monotonically with the translational velocity $U_{\parallel}$ Liu et al. (2024). These observed relationships are a consequence of near-field HIs, which are the physical mechanisms for surface entrapment in swimming bacteria. These interactions result in reductions in both their translational and rotational velocities and in bacterial reorientation. These near-field HIs, along with DLVO forces, collectively determine the stable height $h^{*}$ and the inclination angle $\Psi^{*}$, which subsequently influence the translational velocity, rotational velocity, and radius of curvature of the circular motion. However, the effective range of DLVO forces is significantly shorter, often by several orders of magnitude, than that of near-field HIs. Consequently, the bacterial dynamic behavior on the surface is dominated by near-field HIs.
III.2 Simulation with Brownian Motion
@@
The resistance matrix (Eq. 4) for the cell body near a no-slip solid surface includes three parts: $A_{b}$, which couples force and translational velocity; $C_{b}$, which couples torque and rotational velocity; and $B_{b}$, which couples torque and translational velocity. The submatrices $A_{b}$ and $C_{b}$ indicate that as the bacterium approaches the surface, the hydrodynamic resistance to its translational and rotational motions increases, which leads to a reduction in both translational and rotational velocities. This implies that the bacterium experiences increasingly restricted motion as it approaches the surface. The submatrix $B_{b}$ reveals that shear flow induced by a no-slip solid surface causes a rolling motion of the cell body, which in turn causes the bacteria to orient toward the surface Das et al. (2015); Liu, Chen, and Zhang (2025).
Figure 5: Three different models of bacteria near a surface: (a) Without near-field HIs. (b) Neglecting the submatrix $B_{b}$ of the resistance matrix $\mathcal{R}_{b}$. (c) Including complete near-field HIs.
@@
Figure 6: Sketch diagram of a bacterial trajectory, the bacterium enters and leaves the surface at critical height $h_{c}=7$ $\mu\mathrm{m}$. The surface residence time of bacterium is defined as $\tau$.
@@
In the absence of Brownian motion, bacteria are effectively trapped on the surface under the combination of near-field HIs and DLVO forces, exhibiting stable motion that prevents escape. However, the presence of Brownian motion introduces orientational diffusion, allowing bacteria to eventually escape from the surface Li, Tam, and Tang (2008); Junot et al. (2022). Specifically, near-field HIs suppress both the translational and rotational velocities of bacteria as they approach the surface, potentially prolonging their surface residence time. Furthermore, the non-zero submatrix $B_{b}$ promotes the orientation of bacteria toward the surface, further enhancing their surface residence time. To quantify the influence of near-field HIs on surface residence time, we conduct simulations for three distinct models presented in Fig. 5: Model 1Model\,1, which excludes near-field HIs; Model 2Model\,2, which omits the non-zero submatrix $B_{b}$; and Model 3Model\,3, which incorporates complete near-field HIs.
@@
We define the near-surface region as the area where the closest distance $h$ between the cell body and the surface is less than a critical height $h_{c}=7.0$ $\mu\mathrm{m}$, as shown in Fig. 6. The surface residence time $\tau$ is defined as the duration between the first and last time that the bacteria cross the critical height $h_{c}$ Junot et al. (2022). In subsequent simulations (Fig. 7), the bacteria start with an initial inclination angle of $\Psi=145^{\circ}$ and an initial height of $h=h_{c}$.
@@
Figure 7: Simulations conducted for Debye lengths: $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. (a) Projection of five simulated bacterial trajectories onto the xyxy-plane for $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$. (b) The corresponding heights $h$ as a function of time for these five trajectories. (c) Projection of five simulated bacterial trajectories onto the xyxy-plane for $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. (d) The corresponding heights $h$ as a function of time for these five trajectories. The red dot is the initial position.
@@
Different Debye lengths result in different stable heights $h^{*}$ of bacteria on the surface. To obtain different stable heights, we perform numerical simulations using two characteristic Debye lengths: $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. Fig. 7 shows the projections of five different simulated trajectories onto the surface, along with the evolution of their corresponding heights over time. Figs. 7(a) and (b) present the simulation results for $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$, while Figs. 7(c) and (d) correspond to $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. It can be observed from Fig. 7 that bacteria with left-handed helical flagella tend to swim in clockwise circles. Furthermore, surface residence time increases as stable height $h^{*}$ decreases.
@@
Figs. 7(b) and (d) illustrate the dependence of surface residence time on stable heights. However, the typical range of stable heights we observe, spanning from a few nanometers to more than a hundred nanometers Petroff and Libchaber (2018); Liu, Jin, and An (2025), is significantly smaller than the critical height $h_{c}=7.0$ $\mu\mathrm{m}$. This observation implies that the influence of stable height on surface residence time can ultimately be attributed to near-field HIs. To more precisely quantify the influence of individual components of near-field HIs on surface residence time, we conduct numerical simulations using the parameters provided in Table 1 and the three distinct models depicted in Fig. 5. $10^{4}$ individual bacteria are simulated for each Debye length, and the surface residence time distributions are shown in Figs. 8(a)-(c). The initial height of each bacterium is $h=7.0$ $\mu\mathrm{m}$, and the initial inclination angle is $\Psi=145^{\circ}$.
@@
Figure 8: (a)-(c) Distributions of the surface residence time for Model 1Model\,1, Model 2Model\,2, and Model 3Model\,3, respectively, at Debye lengths of $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$, $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$, and $\kappa^{-1}=6$ $\mathrm{n}\mathrm{m}$. (d) Mean surface residence time $\langle\tau\rangle$ as a function of $\kappa^{-1}$ for these models.
@@
We present the surface residence time distributions for Model 1Model\,1, Model 2Model\,2, and Model 3Model\,3 in Figs. 8(a)-(c), respectively. For Model 1Model\,1 (Fig. 8(a)), decreasing the Debye length results in a slight narrowing of the distribution because the range of inclination angles that capture bacteria on the surface is smaller at $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ than at larger $\kappa^{-1}$ Liu, Jin, and An (2025), resulting in shorter surface residence times. Model 2Model\,2 (Fig. 8(b)), which incorporates submatrices $A_{b}$ and $C_{b}$, exhibits enhanced hydrodynamic resistance at smaller separation distances, leading to a reduced vertical translational velocity of bacteria (Fig. 3). Consequently, the surface residence time distribution at $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ is narrower than those for $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=6$ $\mathrm{n}\mathrm{m}$. Model 3Model\,3 (Fig. 8(c)), incorporating complete near-field HIs, demonstrates that decreasing the stable height results in longer surface residence times. The main distinction between Model 3Model\,3 (Fig. 8(c)) and Model 2Model\,2 (Fig. 8(b)) is the presence of $B_{b}$ in the resistance matrix. $B_{b}$ represents the coupling between translational velocity and torque, which drives the bacteria to orient toward the surface, and thus increases the surface residence time significantly.
@@
We perform numerical simulations involving $10^{4}$ individual bacteria for each model and each Debye length. The mean residence times for each model are illustrated in Fig. 8(d), which intuitively compares the mean residence times across the three models by showing their respective values. It can be observed that the submatrices $A_{b}$ and $C_{b}$ of the resistance matrix cause a slight increase in the mean residence time. In contrast, the submatrix $B_{b}$ significantly increases the residence time, with this effect becoming more pronounced at smaller stable heights where these interactions are amplified. When calculating the mean residence time using Model 1Model\,1, a reduction in Debye length results in an increase in the stable inclination angle, which in turn narrows the range of inclination angles of capturing bacteria on the surface Liu, Jin, and An (2025).
@@
Near-field HIs between bacteria and a no-slip solid surface are the primary physical mechanism underlying surface entrapment, leading to their accumulation near the surface. We employ a chiral two-body model to simulate the dynamics of bacteria near the surface. The surface entrapment process can be divided into three sequential stages: approach, reorientation, and surface swimming. Bacterial motility near the surface is governed by a resistance matrix. This matrix is mainly composed of three types of couplings: $A_{b}$ representing the force-translational velocity coupling; $C_{b}$ representing the torque-rotational velocity coupling; and $B_{b}^{T}$ and $B_{b}$, representing the cross-couplings. As bacteria approach the surface, $A_{b}$ and $C_{b}$ generate an increasing hydrodynamic resistance, resulting in a decrease in vertical translational velocity $U_{\perp}$ and rotational velocity $W_{\perp}$. Upon contact with the surface, $U_{\perp}$ drops to zero, completely confining the bacteria to the surface.
However, the presence of DLVO forces prevents the bacteria from approaching the surface infinitely. Under the combination of near-field HIs and the DLVO force, bacteria reach a stable fixed point $\{h^{*},\Psi^{*}\}$ in the phase plane, where they exhibit circular trajectories on the surface. Notably, bacteria with left-handed helical flagella display clockwise circular motion when viewed from above. During this stage, the horizontal translational velocity $U_{\parallel}$ increases, while the vertical rotational velocity $W_{\perp}$ decreases as the stable height $h^{*}$ increases, collectively resulting in a larger radius of curvature. Additionally, the radius of curvature of the circular motion of bacteria is linearly correlated with their horizontal translational velocity $U_{\parallel}$ for $h^{*}\geq 40$ $\mathrm{n}\mathrm{m}$. Although the submatrices $A_{b}$ and $C_{b}$ have a minor effect on the surface residence time of bacteria, the submatrix $B_{b}$ significantly prolongs this time. In particular, lower stable heights are associated with longer surface residence times. Collectively, these findings suggest that near-field HIs may promote colloid crystallization, a phenomenon that warrants further investigation.
Acknowledgements.

### v2

@@ 3.1 Surface Swimming Dependent on Near-Field HIs
Fig. 3: The normalized velocities of a bacterium approaching the surface perpendicularly. (a) The vertical translational velocity $U_{z}$ as a function of reduced height $h/R_{b}$ normalized by the bulk velocity $U_{\infty}$. The inset is a schematic. (b) The vertical rotational velocity $W_{z}$ as a function of reduced height $h/R_{b}$ normalized by the bulk velocity $W_{\infty}$.
@@
As shown in Fig. 3, we plot the normalized vertical translational velocity of the cell body, $U_{z}/U_{\infty}$, and rotational velocities, $W_{z}/W_{\infty}$, as a function of the reduced height $h/R_{b}$. The translational velocity of bacteria in the bulk fluid is approximately $U_{\infty}=14.5$ $\mu\mathrm{m}\mathrm{/}\mathrm{s}$ for the parameters given in Table 1, which is consistent with the experimental measurements 37. In Figs. 2(a) and (c), the resistance function $Y_{\perp}^{A}$ increases significantly faster than $Y_{\perp}^{C}$ as the separation distance $h$ decreases. Consequently, the first equation of Eq. 14 implies that the vertical translational velocity $U_{z}$ monotonically decreases as the separation distance $h$ decreases, as depicted in Fig. 3(a). The function $Y_{\perp}^{A}$ diverges to infinity as $h$ approaches zero, resulting in the vertical translational velocity of the cell body approaching zero. However, the DLVO force prevents the bacteria from contacting the surface. There is a stable separation distance between the cell body and the surface, which typically ranges from a few nanometers to over a hundred nanometers 41, 32.
Similarly, Fig. 3(b) shows the normalized vertical rotational velocity of the cell body, $W_{z}/W_{\infty}$, as a function of the reduced height $h/R_{b}$. This rotational velocity is also determined by the coupling coefficients $Y_{\perp}^{A}$ and $Y_{\perp}^{C}$, as shown in the second equation of Eq. 14. In the absence of a surface, the bacterial rotational velocity is approximately $W_{\infty}=18.3$ $\mathrm{r}\mathrm{a}\mathrm{d}\mathrm{/}\mathrm{s}$, according to the parameters given in Table 1. Because RFT neglects the hydrodynamic interaction between the flagellum and the cell body, the rotational velocity of the cell body is lower than the experimental measurement 33. However, for qualitatively investigating the velocity change of bacteria approaching the surface, this discrepancy is acceptable. Observing the second equation of Eq. 14, both the numerator and the denominator inherently involve the resistance term $Y_{\perp}^{A}$. As the separation distance $h$ decreases, the vertical rotational velocity of the cell body, $W_{z}$, is jointly influenced by $Y_{\perp}^{A}$ and $Y_{\perp}^{C}$. These functions represent the resistance components for the translational and rotational velocities of the cell body, respectively. The net rotational velocity arises from the competition between these two resistance functions. Consequently, as depicted in Fig. 3(b), the vertical rotational velocity exhibits a slight increase initially before decreasing.
Fig. 4: (a) Stable heights $h^{*}$ and (b) stable inclination angles $\Psi^{*}$ as a function of Debye length for different values of zeta potentials $\zeta$.
During the bacterial surface entrapment process, only hydrodynamic and DLVO interactions are considered. Bacterial surface entrapment is typically divided into three sequential stages: approach, reorientation, and surface swimming 4, 5, 32. Near a no-slip solid surface, a stable fixed point $\{h^{*},\Psi^{*}\}$ exists within the phase plane $\{h,\Psi\}$ 45, 46, where the bacterium performs stable circular motions on the surface. Fig. 4 illustrates the dependence of stable heights ($h^{*}$) and stable inclination angles ($\Psi^{*}$) on the Debye length, with curves shown for varying zeta potentials. These stable states are mainly determined by the bacterial morphology and the Debye length 32.
@@
Fig. 5: Bacterial velocities and radius of curvature of their circular motion in the surface swimming stage. (a) Horizontal translational velocity, $U_{\parallel}$, as a function of stable height $h^{*}$. (b) Vertical rotational velocity, $W_{\perp}$, as a function of stable height $h^{*}$. (c) Radius of curvature, $r$, as a function of stable height $h^{*}$. (d) Radius of curvature, $r$, as a function of horizontal translational velocity $U_{\parallel}$.
@@
Experimental observations have revealed that the radius of curvature is dependent on the bacterial swimming speed 35. Under conditions of fixed bacterial morphology and motor rotation rate, the horizontal translational velocity and vertical rotational velocity of bacteria are solely determined by the stable fixed point $\{h^{*},\Psi^{*}\}$. Bacteria exhibit circular motion near a solid surface during the surface swimming stage 25. The radius of curvature, $r$, of this circular motion is given by 32
@@
where $\Psi^{*}$ is the stable inclination angle of bacteria, $\phi$ is the azimuth angle of the rotational velocity in spherical coordinates. The translational velocity of bacteria parallel to the surface is $U_{\parallel}=\sqrt{U_{x}^{2}+U_{y}^{2}}$, and the rotational velocity of bacteria perpendicular to the surface (along the $z$-axis) is $W_{\perp}=|W_{z}-\left(W_{x}\cos\phi+W_{y}\sin\phi\right)\cot\Psi^{*}|$. The terms $W_{x}$, $W_{y}$, $W_{z}$ are the components of the rotational velocity of the cell body in Cartesian coordinates.
@@
Observations from Eqs. 23-27 reveal that individual resistance functions of the cell body exhibit different dependencies on the separation distance $h$. Figs. 5(a)-(b) show that the horizontal translational velocity $U_{\parallel}$ increases with stable height $h^{*}$, while the vertical rotational velocity $W_{\perp}$ decreases as $h^{*}$ increases. As shown in Fig. 5(c), the radius of curvature of the bacterial circular motion increases monotonically with stable height. For stable heights that exceed $40$ $\mathrm{n}\mathrm{m}$, the radius of curvature exhibits a linear relationship with $h^{*}$ 35. As shown in Fig. 5(d), the radius of curvature also increases monotonically with the translational velocity $U_{\parallel}$ 35. These observed relationships are a consequence of near-field HIs, which are the physical mechanisms for surface entrapment in swimming bacteria. These near-field HIs, along with DLVO forces, collectively determine the stable height $h^{*}$ and the inclination angle $\Psi^{*}$, which subsequently influence the translational velocity, rotational velocity, and radius of curvature of the circular motion. However, the effective range of DLVO forces is significantly shorter, often by several orders of magnitude, than that of near-field HIs. Consequently, the bacterial dynamic behavior near the surface is dominated by near-field HIs.
3.2 Near-Field HIs Prolong Surface Residence Time
@@
The resistance matrix (Eq. 4) for the cell body near a no-slip solid surface comprises three submatrices: $A_{b}$, which describes the coupling between force and translational velocity; $C_{b}$, which describes the coupling between torque and rotational velocity; and $B_{b}$, which describes the coupling between torque and translational velocity. The elements of the submatrices $A_{b}$ and $C_{b}$ reveal that as the bacterium approaches the surface, the hydrodynamic resistance to its translational and rotational motions increases. This consequently leads to a reduction in both translational and rotational velocities, implying increasingly restricted motion for the bacterium near the surface. In contrast, the submatrix $B_{b}$ reveals that shear flow induced by the no-slip solid surface generates a rolling motion of the cell body, which in turn promotes the bacteria to tend to orient towards the surface 11, 31.
@@
Fig. 6: Sketch diagram of a bacterial trajectory, the bacterium enters and leaves the surface at critical height $h_{c}=7$ $\mu\mathrm{m}$. The surface residence time of bacteria is defined as $\tau$.
@@
In the absence of Brownian motion, bacteria are effectively trapped on the surface under the combination of near-field HIs and DLVO forces, exhibiting stable motion on the surface. However, the presence of Brownian motion introduces translational and rotational diffusion, allowing bacteria to eventually escape from the surface 29, 22. However, as bacteria approach the surface, near-field HIs suppress both their translational and rotational diffusion (as shown in Appendix C), potentially prolonging their surface residence time. Furthermore, the non-zero submatrix $B_{b}$ promotes the orientation of bacteria towards the surface, further prolonging their surface residence time.
@@
We define the near-surface region as the area where the closest distance $h$ between the cell body and the surface is less than a critical height $h_{c}=7.0$ $\mu\mathrm{m}$ 22, as illustrated in Fig. 6. The surface residence time $\tau$ is defined as the duration between the first and last instances that the bacteria cross the critical height $h_{c}$ 22. In subsequent simulations (Fig. 7), the bacteria start with an initial inclination angle of $\Psi=145^{\circ}$ and an initial height of $h=h_{c}$.
@@
Fig. 7: Simulations conducted for Debye lengths: $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. (a) Projection of five simulated bacterial trajectories onto the xyxy-plane for $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$. (b) The corresponding heights $h$ as a function of time for these five trajectories. (c) Projection of five simulated bacterial trajectories onto the xyxy-plane for $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. (d) The corresponding heights $h$ as a function of time for these five trajectories. The red dot is the initial position.
@@
Different Debye lengths result in different stable heights $h^{*}$ of bacteria on the surface 32. To obtain different stable heights, we perform numerical simulations using two different Debye lengths: $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. Fig. 7 shows the projections of five different simulated trajectories onto the surface, along with the evolution of their corresponding heights over time. Figs. 7(a) and (b) present the simulation results for $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$, while Figs. 7(c) and (d) correspond to $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$. It can be observed from Fig. 7 that bacteria with left-handed helical flagella tend to swim in clockwise circles. Furthermore, the surface residence time exhibits a corresponding change with the decrease in stable height $h^{*}$.
@@
Fig. 4 shows a strong correlation between stable height and Debye length, and Figs. 7(b) and (d) reveal that surface residence time is closely related to Debye length. However, the quantitative statistical relationship between surface residence time and Debye length remains unclear. Previous studies indicate that different Debye lengths lead to varying stable heights, ranging from a few nanometers to over $100$ nanometers 41, 32, which is significantly smaller than the critical height $h_{c}=7.0$ $\mu\mathrm{m}$. This observation suggests that the influence of stable height on surface residence time can ultimately be attributed to near-field HIs. To quantify the influence of near-field HIs on surface residence time, we perform simulations using the parameters provided in Table 1 and three different models depicted in Fig. 8. Model 1Model\,1, which excludes near-field HIs; Model 2Model\,2, which omits the non-zero submatrix $B_{b}$; and Model 3Model\,3, which incorporates complete near-field HIs. For each model, $10^{4}$ individual bacteria are simulated for each Debye length, and the surface residence time distributions are shown in Figs. 8(a)-(c). The initial height of each bacterium is $h=7.0$ $\mu\mathrm{m}$, and the initial inclination angle is $\Psi=145^{\circ}$.
@@
Fig. 8: (a)-(c) Distributions of the surface residence time for Model 1Model\,1, Model 2Model\,2, and Model 3Model\,3, respectively, at Debye lengths of $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$, $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$, and $\kappa^{-1}=6$ $\mathrm{n}\mathrm{m}$. Three different models of bacteria near a surface: Model 1Model\,1: without near-field HIs; Model 2Model\,2: neglecting the submatrix $B_{b}$ of the resistance matrix $\mathcal{R}_{b}$; Model 3Model\,3: including complete near-field HIs. (d) Mean surface residence time $\langle\tau\rangle$ as a function of $\kappa^{-1}$ for these three models.
@@
In Model 1Model\,1 (Fig. 8(a)), an increase in Debye length leads to a slight broadening of the surface residence time distribution 32. Model 2Model\,2 (Fig. 8(b)), which incorporates submatrices $A_{b}$ and $C_{b}$, exhibits enhanced hydrodynamic resistance at smaller separation distances. This reduction in distance leads to a decrease in the vertical translational velocity of the cell body (Fig. 3). As a result, the surface residence time distribution at $\kappa^{-1}=2$ $\mathrm{n}\mathrm{m}$ is observed to be broader than that for $\kappa^{-1}=4$ $\mathrm{n}\mathrm{m}$ and $\kappa^{-1}=6$ $\mathrm{n}\mathrm{m}$. Incorporating complete near-field HIs, Model 3Model\,3 (Fig. 8(c)) shows that decreasing the Debye length results in broader surface residence time distributions. The main distinction between Model 3Model\,3 (Fig. 8(c)) and Model 2Model\,2 (Fig. 8(b)) is the presence of matrix term $B_{b}$. This term represents the coupling between translational velocity and torque, which drives bacteria to orient toward the surface, thereby significantly prolonging their surface residence times.
@@
For each model and Debye length, numerical simulations are performed using $10^{4}$ individual bacteria to qualitatively determine the average residence time. The mean residence times for each model are illustrated in Fig. 8(d), allowing an intuitive comparison of their values across the three models. It can be observed that the submatrices $A_{b}$ and $C_{b}$ of the resistance matrix cause an increase in the mean residence time. Similarly, the submatrix $B_{b}$ significantly increases the residence time, an effect that becomes more pronounced at smaller stable heights due to the amplification of these specific interactions. When calculating the mean residence time using Model 1Model\,1, a reduced Debye length results in a slight decrease in the mean residence time of bacteria on the surface.
@@
Near-field HIs between bacteria and a no-slip solid surface are the primary physical mechanism underlying surface entrapment, which drives their accumulation near the surface. We employ a chiral two-body model to simulate the dynamics of bacteria near the surface. The surface entrapment process can be divided into three sequential stages: approach, reorientation, and surface swimming. Bacterial motility near the surface is governed by a resistance matrix. This matrix mainly comprises three types of couplings: $A_{b}$ or force-translational velocity coupling, $C_{b}$ for torque-rotational velocity coupling and $B_{b}^{T}$ and $B_{b}$ for cross couplings. As bacteria approach the surface, the influence of $A_{b}$ and $C_{b}$ leads to increased hydrodynamic resistance, causing a decrease in vertical translational velocity $U_{\perp}$. Upon contact with the surface, $U_{\perp}$ becomes zero, effectively confining the bacteria to the surface.
However, the presence of DLVO forces prevents bacteria from contacting the surface. Under the combination of near-field HIs and the DLVO force, bacteria reach a stable fixed point $\{h^{*},\Psi^{*}\}$ in the phase plane, exhibiting circular trajectories near the surface. In particular, bacteria with left-handed helical flagella exhibit clockwise circular motion when viewed from above. During this stage, as the stable height $h^{*}$ increases, the horizontal translational velocity $U_{\parallel}$ increases correspondingly, while the vertical rotational velocity $W_{\perp}$ decreases, which together leads to a larger radius of curvature. Additionally, the radius of curvature of the bacterial circular motion is linearly correlated with the stable height $h^{*}$ for $h^{*}\geq 40$ $\mathrm{n}\mathrm{m}$. Submatrices $A_{b}$ and $C_{b}$ increase the residence time of bacteria on the surface, while the submatrix $B_{b}$ shows a more pronounced effect in prolonging it.
Our research reveals that the near-field HIs between the spherical particles and a solid surface cause the translational diffusion coefficients of the particles to decrease and ultimately reach zero as the particles approach the surface. This significantly prolongs their residence time near the surface. Furthermore, previous work 31 demonstrates that similar near-field HIs also exist between spherical particles. As two spherical particles approach each other, their translational diffusion coefficients decrease to zero, significantly increasing their timescale for separation. Building on these findings and the theory of interparticle near-field HIs, we hypothesize that these interactions promote particle crystallization when the following two conditions are met: 1. The particle size is significantly larger than the effective interaction range of DLVO forces (typically on the micrometer scale), making near-field HIs prominent. 2. The resistance matrix for the system is valid when a no-slip boundary condition is satisfied between the spherical particles. This is the focus of our next research.
Author contributions
Baopi Liu: Data curation, Formal analysis, Methodology, Software, Validation, Visualization, Writing-original draft. Lu Chen: Investigation, Methodology, Validation, Visualization. Haiqin Wang: Investigation, Methodology, Validation, Visualization. All authors reviewed and edited the manuscript.
Conflicts of interest
There are no conflicts of interest to declare.
Acknowledgements